OpenStudy (anonymous):

Determine if the graph is symmetric about the x-axis, the y-axis, or the origin. r = 4 - 3 sin θ @agent0smith i put yaxis only

OpenStudy (agent0smith):

Erm... idk for sure, graph it and find out?

OpenStudy (anonymous):

what about this one? The graph of a limacon curve is given. Without using your graphing calculator, determine which equation is correct for the graph

OpenStudy (anonymous):

[-5, 5] by [-5, 5]

OpenStudy (agent0smith):

^I think i'd need the graph.. if you have a book, it should have some general equation forms.

OpenStudy (agent0smith):

http://mathworld.wolfram.com/Limacon.html what are the options?

OpenStudy (anonymous):

r = 2 + 2 sin θ r = 3 + sin θ r = 1 + 3 cos θ r = 1 + 3 sin θ

OpenStudy (agent0smith):

You could graph each and see which matches, cos i can't tell...